# ETH Zurich

### v1

@@ 3.3 Specializing for single tasks
We now study how much accuracy we stand to gain by fine-tuning on a single task. We specialize models for each of the 10 tasks highlighted in Section 2.4. We specialize the follow models: Llama 3 8B Inst, Llama 3 8B Inst fine-tuned for one epoch on all tasks, and Lawma 8B (i.e., Llama 3 8B Inst fine-tuned for three epochs on all tasks). For each task, we fine-tune for a maximum of 20 epochs and early stop when validation loss increases for 3 consecutive evaluation stpes, each corresponding to one tenth of an epoch.
Figure 9 shows the results of specialization to single tasks. First, we observe that, for 7 out of 10 tasks, Llama 3 8B Inst fine-tuned on all tasks for one epoch (yellow) outperforms Llama 3 8B Inst specialized for a single task (blue). That is, there is value to fine-tuning on our entire dataset rather than overspecializing for a single task. One explanation is that there is substantial cross-task overlap, and fine-tuning on the entire dataset amounts training on many more examples –even if on average these examples are less relevant.
Secondly, we observe that after fine-tuning on all 260 tasks for 1 epoch (yellow), further specializing for a single task (green) improves performance on all cases. Importantly, the latter outperforms the specialized Llama 3 8B Inst (blue) in all tasks. That is, a model that is fine-tuned on everything provides a “better” foundation from which to then “overspecializing” for a single task.
Thirdly fine-tuning on everything for three epochs (i.e., Lawma 8B, in red) again improves over the specialized models (i.e., green). Lastly, “overspecializing” Lawma 8B for a single task results in small single digit improvements for 3 out of the 10 tasks. However, we observe no benefits from specializing Lawma 8B for most (7/10) of the tasks. [^4] There is a small decrease in performance for SC Issue Area. This is because early stopping is performed with respect to loss on the validation set, but models are evaluated for accuracy on the test set. These results show that we don’t leave much accuracy on the table by fine-tuning a single model for all tasks. This is practically quite appealing, since it obviates the need to maintain a separate model for each task. A single model suffices.
@@
3.4 Generalization to unseen databases
We now investigate whether fine-tuning only on the Songer Appeals Court database allows us to generalize to the Supreme Court database. We fine-tune Llama 3 8B Inst for one epoch on all Songer tasks simultaneously. We plot in Figure 10 the mean accuracy for Court of Appeals tasks and Supreme Court tasks at intermediate checkpoints. As expected, performance on Court of Appeals tasks improves monotonically with the number of training examples seen. More interestingly, we observe that mean task accuracy for the Supreme Court also improves substantially, by up to 18.8 accuracy points at 20% of the training steps [^5] Note that 20% is the optimal amount of Songer data to train on if the goal is to generalize precisely to Supreme Court cases. If the goal is to generalize to some other dataset, then 20% need not be the optimal amount of Songer data to train on.. Thereafter, performance degrades, seemingly plateauing at 11.3 accuracy points above the baseline non-finetuned performance of Llama 3 8B Inst.
Our findings indicate that, since there is some degree of overlap between Court of Appeal and Supreme Court tasks, fine-tuning on the former transfers to the latter. This suggests that Lawma might be of practical use beyond the Supreme Court and Court of Appeals tasks it was trained on.
Note, however, that fine-tuning only on the Court of Appeals database results in a mean case accuracy of 51.6%, compared to 82.4% for Lawma 8B. That is, not fine-tuning on Supreme Court cases results in a 30.9 accuracy points decrease in performance. These results again highlight the importance of fine-tuning precisely on the target tasks of interest.
Figure 10: Fine-tuning only on the Court of Appeals database improves accuracy on Supreme Court database.
@@ 3.5 Intercoder agreement analysis
Table 2 considers several tasks from the Appeals Court database, including the selected ones we highlighted in various figures. Each row corresponds to one task and provides the intercoder agreement rate, adjusted (and unadjusted) accuracy achieved by Lawma 8B, and the fraction of samples we retained in the majority class. A fraction of 100% means that we kept all samples. The smaller the fraction the larger the majority class is relative to the other classes.
The table contains several interesting insights:
@@
These findings speak to the task heterogeneity and the non-trivial nature of the task suite as a classification benchmark.
@@ 4 Discussion
The few-shot capabilities of large language models are vital for commercial APIs, where users are largely restricted to prompting. But as we show, zero-shot prompting is neither sufficiently good nor best possible for classification tasks that arise in empirical legal work. Lightly fine-tuned special purpose models achieve significantly higher accuracy from relatively few labeled examples. Labeling a few hundred cases is often financially feasible. This suggests a simple and practical strategy for solving legal classification tasks: Obtain a few hundred labeled examples, fine-tune an open source model, and use the fine-tuned model to annotate the remaining cases.
Multi-shot prompting is unlikely to provide much relief about the limitations of zero-shot prompting. Many legal documents are so long that a single document essentially consumes the entire prompt window. For example, 38% of court opinions in our dataset have more than 4,000 tokens. The default GPT-4 context size is 8,000 tokens. A version of GPT-4 allowing for 32,000 tokens is substantially more expensive. Fine-tuning avoids this problem altogether and robustly handles long documents.
@@ Acknowledgments
Appendix A Fine-tuning details
@@
A.1 Lawma
We fine-tuning with a maximum sequence length of 8192 tokens. We use the AdamW optimizer with full precision, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$. We use a peak learning rate of $2\cdot 10^{-6}$. We use a cosine learning rate schedule, with 180 warm-up steps (approx. 4% of a full epoch) and decay to $10\%$ of the peak learning rate. We use a weight decay of 0.1. We clip gradient to 1.0 max norm. We pack samples using the axolotl library (Cloud, 2024), which improves training efficiency by approximately 40%. For Lawma 8B, we fine-tune Llama 3 8B Instruct for 3 epochs. We train on a node of 7 H100s using DeepSpeed Zero 2, with a global batch size of 56. For Lawma 70B, we fine-tune Llama 3 70B Instruct for 1 epoch. We train on 8 nodes of 8 H100s each using DeepSpeed Zero 3, with a global batch size of 64. We find that additional epochs hurt average task performance, although performance continues to improve for some of the tasks.
A.2 Additional fine-tuning experiments
@@
Appendix B List of all tasks

### v2

@@ 4.4 Specializing for single tasks
We now investigate how much accuracy can be gained by fine-tuning on a single task. To this end, we specialize models for each of the 10 highlighted annotation tasks. Specifically, we evaluate the following models: Llama 3 8B Instruct, Llama 3 8B Instruct fine-tuned for one epoch on all tasks, and Lawma 8B (i.e., fine-tuned for three epochs on all tasks). For each individual task, we fine-tune for up to 20 epochs, early stopping when the validation loss increases for three consecutive tenths of an epoch.
@@
Figure 13 shows the results of task-specific fine-tuning. First, we observe that for 7 out of the 10 tasks, the model fine-tuned on all tasks for one epoch (yellow) outperforms the model specialized on a single task (blue). This suggests that fine-tuning on the entire dataset provides more benefit than only specializing for a single task. A plausible explanation is that there is substantial overlap among tasks; thus, fine-tuning on the entire dataset exposes the model to many more training examples—even if some are less directly relevant.
Second, we find that starting from a model already fine-tuned on all 260 tasks for one epoch (yellow), then further specializing it on a single task (green), leads to performance improvements across all tasks. Notably, this approach consistently outperforms the models that were specialized on a single task from scratch (blue). In other words, a broadly fine-tuned model serves as a better foundation for subsequent specialization.
Third, fine-tuning on all tasks for three epochs (i.e., Lawma 8B, red) yields additional improvements over both the yellow and green models. Finally, further specializing Lawma 8B on individual tasks leads to modest, single-digit accuracy gains in 3 of tasks. However, for the remaining 7 tasks, no further benefit is observed. [^1] For SC Issue Area, performance slightly decreases. This is likely due to a mismatch between the early stopping criterion (validation loss) and the evaluation metric (test accuracy).
These results indicate that we don’t leave much accuracy on the table by fine-tuning a single model for all tasks. This is practically appealing, since it obviates the need to maintain a separate model for each task.
@@ 4.5 Intercoder agreement analysis
Figure 12 shows the adjusted accuracies of Lawma 8B plotted against intercoder agreement. We observe substantial variation in how model accuracy correlates with intercoder agreement. Notably, for many tasks, the model achieves only moderate performance (e.g., 80–90%) even when intercoder agreement is high (e.g., above 95%). On average across Courts of Appeals tasks, model performance lags behind intercoder agreement by 7.6 percentage points, indicating significant room for improvement.
We highlight several specific tasks in Table 1. Each row corresponds to one task and provides the intercoder agreement rate, adjusted (and unadjusted) accuracy achieved by Lawma 8B, and the fraction of samples we retained in the majority class. A fraction of 100% means that we kept all samples. The smaller the fraction the larger the majority class is relative to the other classes. The table contains several interesting insights:
@@
These findings speak to the task heterogeneity and the non-trivial nature of the CaselawQA suite of legal annotation tasks as a classification benchmark.
@@ 5 Discussion
The generalist abilities of large language models are vital for commercial APIs, where users are largely restricted to prompting. But as we show, generalist models are neither sufficiently good nor best possible for annotation tasks that arise in empirical legal work. Lightly fine-tuned special purpose models achieve significantly higher accuracy from relatively few labeled examples. Labeling a few hundred cases is often financially feasible. This suggests a simple and highly scalable strategy for solving legal classification tasks: Obtain a few hundred labeled examples, fine-tune an open-weights model, and use the fine-tuned model to annotate the remaining cases.
@@ Acknowledgments
Appendix A Highlighted tasks
Four tasks from the USCAD and all tasks from the SCDB were selected to form pairs, with each pair consisting of one task from the USCAD and one from the SCDB that capture similar concepts. It is important to note that, despite capturing broadly similar concepts, the precise formulation of the tasks might differ between the USCAD and the SCDB, making them less than perfectly comparable. In addition to the four pairs, we include two tasks from the SCDB that involve determining features of the decision reviewed by the Supreme Court on the basis of the Supreme Court opinion. The following is a description of the task pairs:
Appendix B Additional performance results
We present additional performance results with different performance metrics (i.e., balanced accuracy, macro-F1), without subsampling the majority class (i.e., allowing the majority class to account for more than half of the examples), and when computing mean task accuracy (i.e., macro-averaging rather than micro-averaging across tasks). Since we must evaluate on CaselawQA’s extended test set, which contains 14x more examples, we restrict our evaluation to Llama 3 8B Instruct, Llama 3 70B Instruct, GPT-4, and Lawma 8B. We evaluate models using direct question-answering.
Balanced accuracy and macro-F1
See Figure 14 and Figure 15 for evaluation results using mean balanced accuracy and mean macro-F1 as the evaluation metric, respectively.
Figure 14: Evaluation results using balanced accuracy as the evaluation metric.
Figure 15: Evaluation results using mean macro-F1 as the evaluation metric.
Results without subsampling the majority class
Figure 16 presents the evaluation results when not subsampling the majority class. Models achieve very hight accuracy on many tasks simply because they correctly identify the majority class.
Figure 16: Evaluation results without subsampling the majority class.
Appendix C Fine-tuning details
@@
C.1 Lawma
We fine-tune with a maximum sequence length of 8192 tokens. We use the AdamW optimizer with full precision, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$. We use a peak learning rate of $2\cdot 10^{-6}$. We use a cosine learning rate schedule, with 180 warm-up steps (approx. 4% of a full epoch) and decay to $10\%$ of the peak learning rate. We use a weight decay of 0.1. We clip gradient to 1.0 max norm. We pack samples using the axolotl library (Cloud, 2024), which improves training efficiency by approximately 40%. For Lawma 8B, we fine-tune Llama 3.1 8B Instruct for 3 epochs. We train on a node of 7 H100s using DeepSpeed Zero 2, with a global batch size of 56. For Lawma 70B, we fine-tune Llama 3.3 70B Instruct for 3 epochs. We train on 8 nodes of 8 H100s each using DeepSpeed Zero 3, with a global batch size of 64. We find that additional epochs do not significantly improve performance.
C.2 Additional fine-tuning experiments
@@
Appendix D List of all tasks